# Integrating Reinforcement Learning and Model Predictive Control with Applications to Microgrids

### v1

@@ 4 Method: Novel integration of reinforcement learning and MPC
In this section, we present the main contribution of our work: an integrated reinforcement learning and MPC method for the solution of mixed-integer linear and quadratic problems for the control of mixed-logical dynamical systems. The main goal is to ease the online computational burden of solving mixed-integer linear and quadratic programs by decoupling the computation of the integer and continuous decision variables. In our framework, the integer decision variables are determined by reinforcement learning, and the continuous decision variables by MPC, which is an optimization-based framework. Accordingly, the solution of mixed-integer programs is entirely avoided in on-line operation. Consequently, our approach can be seen as an alternative to branch-and-bound for MPC for MLD systems. In what follows, the decoupling of the discrete and continuous variables of the MILP of (5) is addressed. Next, we explain the role of reinforcement in our approach.
Herein it is assumed that the system dynamics are known and training takes place fully offline, i.e., there is no online adaptation of the learning parameters.
@@ 4.1 Decoupling of decision variables
The mixed-integer linear program (MILP) (5) is parametric with respect to the initial state $x(k)$ and possibly an exogenous signal $\gamma(k)$. Consider the set of these values in which (5) is feasible. In this set, the solution of (5) maps the signal $(x(k),\ \gamma(k))$ to their corresponding optimal decision variables $(\epsilon_{\mathrm{c}}(k)^{*},\ \epsilon_{\mathrm{d}}(k)^{*})$. Note that the uniqueness of the solution of (5) is not assumed and multiple optima may exist. In this case, only one solution is considered. Henceforth, we use the augmented state
$\chi(k)=[x^{\mathrm{T}}(k),\ \gamma^{\mathrm{T}}(k)]^{\mathrm{T}}$
to refer to the augmented state $\chi(k)$ containing the system state $x(k)$ and the exogenous signal $\gamma(k)$ at time step $k$.
@@
Consider the scenario where the discrete part of the optimal solution $\epsilon_{\mathrm{d}}(k)^{*}$ is known. Then the continuous part of the solution $\epsilon_{\mathrm{c}}(k)^{*}$ can be easily recovered by solving the problem (5), which becomes a linear program (LP) with the fixed discrete optimization variables. The principle used in our approach to decouple the discrete and continuous decision variables is based on this fact. In the proposed approach, the discrete variables are determined by reinforcement learning. Then solving for the continuous MILP (5) can be simplified into the following LP:
$\begin{split}V(\chi(k),\ \epsilon_{\mathrm{d}}(k))=\min_{\epsilon_{\mathrm{c}}(k)}\ &c_{\mathrm{c}}^{\mathrm{T}}\epsilon_{\mathrm{c}}(k)\\ \text{subj. to }&G[\epsilon_{\mathrm{c}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}\leq w+S\chi(k)\end{split}$ (6)
@@
The main difficulty of solving MILPs is the discrete optimization variables and modern solvers mostly rely on branch-and-bound as a strategy to address this problem. Therefore, computing these complicating decision variables with learning-based methods removes the main optimization hurdle. Furthermore, solving such LPs of the form (6) is computationally cheap with state-of-the-art specific-purpose solvers. The immediate benefit of the continuous variables being determined by an optimization-based control, such as MPC, is constraint satisfaction and optimality – given the feasibility of (6). A depiction of the decoupling process is shown in Figure 1.
@@
Figure 1: Representation of the decoupling of the discrete and continuous decision variables. From mixed logical dynamical (MLD) modeling and the use of an MPC approach for control, a mixed-integer linear program (MILP) can be formulated for the operation of the microgrid. Then, a learning approach – either reinforcement learning (RL) or supervised learning (SL) – is used to determine the discrete variable $\epsilon_{\mathrm{d}}(k)$. The MILP is then simplified into a linear program (LP), which computes the continuous variable $\epsilon_{\mathrm{c}}(k)$.
@@ 4.2 Role of reinforcement learning
In the mixed-integer linear program (5), the main computational complexity stems from the number of discrete decision variables, which can come from both the number of actions per time step and the length of the prediction horizon. In our method, reinforcement learning is used to ease the online computational burden of solving mixed-integer linear programs by decoupling the computation of the discrete and continuous decision variables. In our framework, the discrete decision variables are determined by reinforcement learning, and the continuous decision variables by optimization-based control – MPC. Accordingly, the solution of mixed-integer programs is entirely avoided in on-line operation.
@@
In order to formulate the problem as a Markov decision process (MDP) – the standard reinforcement learning framework – we lump together the system and the MPC controller in a single block to form the environment, see Figure 2. The environment receives the discrete actions from the RL agent and outputs the next state and the corresponding reward. This abstraction allows the decoupling of the discrete and continuous decision variables and it is crucial to bridge reinforcement learning and optimization-based control for MLD systems. The agent is responsible for determining the discrete sequence of actions $\epsilon_{\mathrm{d}}(k)$ over the prediction horizon. This vector is then sent to the environment which solves the optimization problem (6) and computes the continuous decision variables $\epsilon_{\mathrm{c}}(k)$. As a result, the environment outputs the next state $\chi(k+1)$ and the reward $r(k)$. Any reinforcement learning algorithm can be employed to train the agent; however, for simplicity, we use Deep Q-Learning [29] to present our approach. For a graphical representation of the reinforcement learning setting, see Figure 2.
@@
Figure 2: A depiction of the proposed control scheme that integrates reinforcement learning into an MPC framework. The agent’s goal is to maximize its long-term reward. It learns to adapt its policy by repeatedly interacting with the environment, that is, by sending a discrete action $\epsilon_{\mathrm{d}}(k)$ and by receiving the extended state $\chi$ and immediate reward $r$. The MPC controller, which is lumped in the environment, receives this discrete action $\epsilon_{\mathrm{d}}(k)$ and then solves an optimization problem to determine the continuous action $\epsilon_{\mathrm{c}}(k)$. Finally, the input $\epsilon$ is fed to the system and the next state is computed.
@@
Herein, the Q-function is defined as the expected reward over a finite horizon if the agent takes action $\epsilon_{\mathrm{d}}(k)$ at the current time step $k$ and then follows the policy $\epsilon_{\mathrm{d}}(k+l)=\pi(\chi(k+l))$ over the remaining steps of the horizon:
$Q^{\pi}(\chi(k),\epsilon_{\mathrm{d}}(k))=r(k)+\sum_{l=1}^{N_{\mathrm{p}}-1}\alpha^{l}r(k+l)$ (7)
@@
$r(k)=f(V(\chi(k),\ \epsilon_{\mathrm{d}}(k)))$
@@
The vector $\epsilon_{\mathrm{d}}(k)$ contains the entire sequence of discrete variables over the prediction horizon. For each time step of the prediction horizon $l=0,\ldots,N_{\mathrm{p}}-1$, we can represent the discrete action per time step with $\varepsilon_{\mathrm{d},l}(k)$, hence
@@
For example, suppose that all of the discrete optimization variables are binary and the parameter $N_{\mathrm{b}}$ represents the number of binary actions per time step, then there are $2^{N_{\mathrm{b}}\cdot N_{\mathrm{p}}}$ possible sequences for the vector $\epsilon_{\mathrm{d}}(k)$. As the number of binary variables $N_{\mathrm{b}}$ and the length of the prediction horizon $N_{\mathrm{p}}$ increase, the number of sequences grows exponentially. Due to the potentially large size of the action space, this is a very challenging problem for RL algorithms. For instance, efficient exploration becomes a big issue in this setting.
Instead of using an RL algorithm to find the Q-function for the action $\epsilon_{\mathrm{d}}(k)$, as defined in (7), our approach shifts the goal to learning the so-called decoupled Q-functions for the sub-actions $\{\varepsilon_{\mathrm{d},l}(k)\}_{l=0}^{N_{\mathrm{p}}-1}$. The decoupled Q-functions are pseudo Q-functions in the sense that they do not entirely fit the standard definition of a Q-function. These decoupled Q-functions are just used as proxies to determine the sub-actions, i.e.,
$\varepsilon_{\mathrm{d},l}(k)=\max_{\varepsilon}\ Q_{l}(\chi(k),\ \varepsilon,\ h_{l-1})\ {\text{for }l=1,\ \ldots,\ N_{\mathrm{p}}-1}.$ (8)
where $h_{l-1}$ is a signal that depends on the previous actions up to time step $i$ of the prediction horizon $\{\varepsilon_{\mathrm{d},j}(k)\}_{j=0}^{j=i}$. If the sub-actions were independent, the learning problem would be greatly simplified. However, this assumption is very constraining for most useful applications where optimal planning requires knowledge of past sub-actions. While each step of the prediction horizon has its own decoupled Q-function $Q_{l}(\cdot)$, the signal $h_{l-1}$ retains some of the couplings with the past sub-actions. In this way, some inter-dependency between the sub-actions is preserved and the learning problem is simplified with the decoupling of the Q-functions. Note that the size of the action space for each decoupled Q-function is $2^{N_{\mathrm{b}}}$, showing that the combinatorial growth is mitigated with respect to the prediction horizon.
To approximate the decoupled Q-functions, we propose the use of recurrent neural networks (RNNs), which have internal states that carry over information from the past – similar to the role of the signal $h$ in (8). More specifically, a long short-term memory (LSTM) network is employed for its capacity to capture long-term time dependencies in data [30]. It is important to stress that the LSTM is a particular implementation of (8) and similar RNNs can also be used to attain the same objective. An unrolled LSTM is shown in Figure 3 to represent the approximation of the decoupled Q-values. The idea of using an LSTM network to decouple the actions of each time step of the prediction horizon was first explored in [11] in the context of supervised learning. Our approach, however, employs the LSTM network as a function approximator of the decoupled Q-functions in the setting of reinforcement learning.
@@
It is worth stressing that the decision-making between time steps is not fully decoupled because the internal state of the LSTM network carries information forward in time, maintaining the dependence of future actions on past actions. Moreover, the integer and continuous decision variables are also not entirely decoupled because the learning agent is trained based on the augmented state and reward signals, which depend on all the decision variables. Therefore, in both cases, decoupling aids tractability while maintaining some degree of interconnection, which is arguably desirable in our setting.
@@
Figure 3: A representation of the recurrent LSTM network on the left-hand side and the unrolled LSTM network on the right-hand side. Note that the LSTM is unrolled for the duration of the prediction horizon $N_{\mathrm{p}}$. Moreover, at each time step $k$, the augmented state $\chi(k)$ can go through the operation $g_{k}(\cdot)$, changing the manner in which the augmented state is presented at the LSTM network at time step $k$. This can be interpreted as a pre-processing technique to better exploit the structure of our problem.
@@
Each episode of the training procedure starts with a randomly chosen initial state $\chi(k)$. Then, the RL agent repeatedly proposes a discrete action $\epsilon_{\mathrm{d}}(k)$, sends it to the environment, and collects the reward $r(k)$ and the next state $\chi(k+1)$. An episode ends when a pre-determined time limit is reached or when the RL agent suggests a discrete action that makes the optimization problem (6) infeasible. At each time step, the transitions $(\chi(k),\ \epsilon_{\mathrm{d}}(k),\ r(k),\ \chi(k+1))$ are stored in a data buffer $\mathcal{D}$, which only holds a finite number of most recent transitions. After the end of each episode, $T$ transitions are randomly sampled from $\mathcal{D}$ and the weights of the LSTM network are updated by gradient descent on the following loss function:
$L=\frac{1}{T}\sum_{k=1}^{T}\sum_{l=0}^{N_{\mathrm{p}}-1}(y(k)-Q_{l}(\chi(k),\ \varepsilon_{\mathrm{d},l}(k)))^{2}$
where the temporal-difference (TD) target is defined as
$y(k)=r(k)+\alpha\frac{1}{N_{\mathrm{p}}}\sum_{l=0}^{N_{\mathrm{p}}-1}\max_{\varepsilon}Q_{l}(\chi(k),\ \varepsilon)$
and the dependency of the decoupled Q-functions on $h$ has been omitted for the sake of simplicity. The summation on the right-hand side represents the average of the maximum decoupled Q-values for each of the sub-actions. Note that a single TD target is used for all decoupled Q-functions. Some alternative definitions of the TD target can be found in [13]. Training ends when a pre-determined number of episodes or number of transitions is reached.

### v2

@@ 4 METHOD: NOVEL INTEGRATION OF REINFORCEMENT LEARNING AND MPC
In this section, we present the main contribution of our work: an integrated reinforcement learning and MPC method for the solution of mixed-integer linear problems for control of mixed-logical dynamical systems. The main goal is to ease the online computational burden of solving mixed-integer linear programs by decoupling the computation of the integer and continuous decision variables. In our framework, the discrete decision variables are determined by reinforcement learning, and the continuous decision variables by MPC. Accordingly, the solution of mixed-integer programs is entirely avoided in online operation. Consequently, our approach can be seen as an alternative to branch-and-bound for MPC for MLD systems. In what follows, the decoupling of the discrete and continuous variables of the MILP of (6) is addressed. Next, we explain the role of reinforcement in our approach.
Herein, it is assumed that the system dynamics are known and training takes place fully offline, i.e., there is no online adaptation of the learning parameters.
@@ 4.1 Decoupling of decision variables
The mixed-integer linear program (MILP) (6) is parametric with respect to the initial state $x(k)$ and possibly an exogenous signal $\gamma(k)$. Consider the set of these values in which (6) is feasible. In this set, the solution of (6) maps the signal $(x(k),\ \gamma(k))$ to their corresponding optimal decision variables $(\epsilon_{\mathrm{c}}(k)^{*},\ \epsilon_{\mathrm{d}}(k)^{*})$.
@@
Consider the scenario where the discrete part of the optimal solution $\epsilon_{\mathrm{d}}(k)^{*}$ is known. Then the continuous part of the solution $\epsilon_{\mathrm{c}}(k)^{*}$ can be easily recovered by solving the problem (6), which becomes a linear program (LP) with the fixed discrete optimization variables, as shown below:
$\begin{split}V(x(k),\ \gamma(k),\ &\epsilon_{\mathrm{d}}(k)^{*})=\min_{\epsilon_{\mathrm{c}}(k)}\ c(\gamma(k))^{\mathrm{T}}[\epsilon_{\mathrm{c}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{d}}^{\mathrm{T}}(k)^{*}]^{\mathrm{T}}\\ \text{s.t. }&G[\epsilon_{\mathrm{c}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{d}}^{\mathrm{T}}(k)^{*}]^{\mathrm{T}}\leq w(\gamma(k))+Sx(k)\end{split}$ (7)
In this manner, the optimal continuous decision variables can be found, provided that the optimal discrete decision variables are known. Evidently, assuming the knowledge of the discrete decision variables is not realistic. Nevertheless, we make use of this decoupling procedure by using learning approaches to provide approximate solutions to the discrete decision variables and by letting the optimization determine the continuous variables, as in (7). One of the main goals of this work is to show that, with sufficient training data, an effective approximator for the discrete decision variables can be trained and that solving (7) with such approximated discrete variables yields solutions for the continuous variables without much sacrifice of optimality.
In the proposed approach, reinforcement learning is used to learn a policy that maps the system state to the discrete decision variables. The training of RL policy is done offline by repeated interaction with a model of the system and historical data. During online operation, the original MILP problem (6) is turned into an LP problem (7) after the RL policy provides an approximate solution to the discrete decision variables given the current measured system state and exogenous variable. A more detailed description of the role of RL is given in the next subsection.
@@
The main difficulty of solving MILPs is the discrete optimization variables, and modern solvers mostly rely on branch-and-bound as a strategy to address this problem. Therefore, computing these complicating decision variables with a learning-based method removes the main optimization hurdle. Furthermore, solving such LPs of the form (7) is computationally cheap with state-of-the-art specific-purpose solvers. The immediate benefit of the continuous variables being determined by an optimization-based control, such as MPC, is constraint satisfaction and optimality relative to the approximate discrete variables solution and subject to the feasibility of (7). A depiction of the decoupling process is shown in Fig. 1.
@@
Figure 1: Representation of the decoupling of the discrete and continuous decision variables. From mixed logical dynamical (MLD) modeling and the use of an MPC approach for control, a mixed-integer linear program (MILP) can be formulated for the operation of the microgrid. Then, a learning approach—either reinforcement learning (RL) or supervised learning (SL) – is used to determine the discrete variable $\epsilon_{\mathrm{d}}(k)$. The MILP is then simplified into a linear program (LP), which computes the continuous variable $\epsilon_{\mathrm{c}}(k)$.
@@ 4.2 Role of reinforcement learning
In the mixed-integer linear program (6), the main computational complexity stems from the number of discrete decision variables, which can come from both the number of actions per time step and the length of the prediction horizon. In our method, reinforcement learning is used to ease the online computational burden of solving mixed-integer linear programs by decoupling the computation of the discrete and continuous decision variables. The discrete decision variables are determined by reinforcement learning, and the continuous decision variables by optimization-based control – MPC. Accordingly, the solution of mixed-integer programs is entirely avoided in online operation.
@@
In order to formulate the problem as a Markov decision process (MDP) – the standard reinforcement learning framework – we lump together the system and the MPC controller in a single block to form the environment, see Fig. 2. The environment receives the discrete actions $\epsilon_{\mathrm{d}}(k)$ from the RL agent and outputs the next state $\chi(k+1)$ and the corresponding reward $r(k)$. This abstraction allows the decoupling of the discrete and continuous decision variables, and it is crucial to bridge reinforcement learning and optimization-based control for MLD systems. The agent is responsible for determining the discrete sequence of actions $\epsilon_{\mathrm{d}}(k)$ over the prediction horizon. This vector is then sent to the environment, which solves the optimization problem (7) and computes the continuous decision variables $\epsilon_{\mathrm{c}}(k)$. As a result, the environment outputs the next state $\chi(k+1)$ and the reward $r(k)$. Any reinforcement learning algorithm can be employed to train the agent; however, for simplicity, we use Deep Q-Learning [34] to present our approach. For a graphical representation of the reinforcement learning setting, see Fig. 2.
@@
Figure 2: A depiction of the proposed control scheme that integrates reinforcement learning into an MPC framework. The agent’s goal is to maximize its long-term reward. It learns to adapt its policy by repeatedly interacting with the environment, that is, by sending a discrete action $\epsilon_{\mathrm{d}}(k)$ and by receiving the extended state $\chi$ and immediate reward $r$. The MPC controller, which is lumped in the environment, receives this discrete action $\epsilon_{\mathrm{d}}(k)$ and then solves an optimization problem to determine the continuous action $\epsilon_{\mathrm{c}}(k)$. Finally, the input $\epsilon$ is fed to the system, and the next state is computed.
@@
Herein, the Q-function is defined as the expected reward over a finite horizon if the agent takes action $\epsilon_{\mathrm{d}}(k)$ at the current time step $k$ and then follows the policy $\epsilon_{\mathrm{d}}(k+l)=\pi(\epsilon_{\mathrm{d}}\ |\ \chi(k+l))$ over the remaining steps of the horizon:
$\begin{split}Q^{\pi}(\chi,\ &\epsilon_{\mathrm{d}})=\ \mathbb{E}_{\pi}[\sum_{l=0}^{N_{\mathrm{p}}-1}\alpha^{l}r(k+l)\ |\ \chi(k)=\chi,\\ &\epsilon_{\mathrm{d}}(k)=\epsilon_{\mathrm{d}},\ \epsilon_{\mathrm{d}}(k+l)=\pi(\epsilon_{\mathrm{d}}\ |\ \chi(k+l))\ \forall l\neq 0]\end{split}$ (8)
@@
$r(k)=f_{\mathrm{reward}}(\gamma(k),\epsilon(k))$
@@
The vector $\epsilon_{\mathrm{d}}(k)$ contains the entire sequence of discrete variables over the prediction horizon. For each time step of the prediction horizon $l=0,\ldots,N_{\mathrm{p}}-1$, we can represent the discrete sub-action per time step with $\varepsilon_{\mathrm{d},l}(k)$, hence
@@
The potentially large size a combinatorial action space is a very challenging problem for RL algorithms. For instance, estimating the Q-function via neural networks, as it is typically implemented in value-based and actor-critic methods, would require a large and intractable number of units – equal to the number of actions – in the output layer. Instead of using an RL algorithm to find the Q-function for the action $\epsilon_{\mathrm{d}}(k)$, as defined in (8), our approach shifts the goal to learning the decoupled Q-functions for the sub-actions $\{\varepsilon_{\mathrm{d},l}(k)\}_{l=0}^{N_{\mathrm{p}}-1}$. Formally, the decoupled Q-function under a policy $\pi(\epsilon_{\mathrm{d}}\ |\ \chi(k))$ for a given time step $l\in\{0,...,N_{\mathrm{p}}-1\}$ of the prediction horizon is defined as
[…]
where $[\pi(\epsilon_{\mathrm{d}}\ |\ \chi(i))]_{j}$ is the $j$th element of the discrete decision variables vector. More specifically, the decoupled Q-function is defined as the expected return of sub-action $\varepsilon$ in state $\chi$ given that the first sub-action at index $l$ is $\varepsilon_{d,l}(0)=\varepsilon$, the initial state is $\chi(0)=\chi$, the other sub-actions $j\neq l$ are chosen under policy $\varepsilon_{\mathrm{d},j}(0)=[\pi(\epsilon_{\mathrm{d}}\ |\ \chi(k))]_{j}$ for the first time step ($k=0$) and that the system is guided by policy $\epsilon_{\mathrm{d}}(k)=\pi(\epsilon_{\mathrm{d}}\ |\ \chi(k))$ for $k\in\{1,...,N_{\mathrm{p}}-1\}$ under the MDP dynamics $\chi(k+1)=f_{\mathrm{MDP}}(\chi(k),\ \epsilon(k))$. The decoupled Q-functions can then be used to determine the policies for each of the sub-actions, e.g., the greedy policies are defined as
$\varepsilon_{\mathrm{d},l}(k)=\max_{\varepsilon}\ Q_{l}^{\pi}(\chi(k),\ \varepsilon)\ {\text{for }l=0,\ \ldots,\ N_{\mathrm{p}}-1}.$ (10)
The definition (9) is not tractable, but it can be useful for the design of a suitable approximator. The goal is to approximate each of the decoupled Q-functions $\{\tilde{Q}_{l}(\cdot)\}_{l=0}^{N_{\mathrm{p}}-1}$ over the prediction horizon. By definition, the decoupled Q-function of time step $j$ depends on the sub-actions $l\in\{0,...,N_{\mathrm{p}}-1\}\setminus\{j\}$ taken over the prediction horizon via the policy $\pi(\cdot)$. Herein, we assume that the past sub-actions are more relevant to the current sub-action than the future sub-actions. Therefore, to compute the decoupled Q-function of the time step $j$, only the sub-actions up to time step $j$ of the prediction horizon $\{\varepsilon_{d,l}(k)\}_{l=0}^{j-1}$ are considered. We propose the use of recurrent neural networks (RNNs) to recursively compute each of the decoupled Q-values for its capacity of carrying over information from the past through an internal state. In this fashion, the decoupled Q-function receives information from the past sub-actions. Based on this assumption, an approximation of the decoupled Q-function with an RNN can be formulated as
$Q_{l}(\chi(k),\ \varepsilon_{\mathrm{d,l}}(k))\approx\tilde{Q}_{l}(\chi(k),\ \varepsilon_{\mathrm{d,l}}(k),\ h_{l-1}(k))$ (11)
where the hidden state $h_{l-1}(k)$ condenses information about past sub-actions. The evaluation of each decoupled Q-function is illustrated in Fig. 3, which depicts the forward pass of the unrolled long short-term memory (LSTM) network, a type of RNN. In particular, the LSTM network is employed for its capacity of capturing long-term dependencies in data and its structure, which attenuates the vanishing gradient problem [35]. Notably, the online evaluation of this approximator involves only the forward pass of the LSTM network, ensuring computational efficiency and tractability.
@@
While each step of the prediction horizon has its own decoupled Q-function $Q_{l}(\cdot)$, the signal $h_{l-1}(k)$ retains the dependence of the previous sub-actions. In this way, some degree of interdependence between the sub-actions is modeled, and the learning problem is simplified due to the decoupling of the Q-functions. Moreover, the integer and continuous decision variables are also not entirely decoupled since the training of the approximator is based on the augmented state and reward signals, which depend on all the decision variables. Therefore, in both cases, decoupling aids tractability while maintaining some degree of interconnection, which is arguably desirable in our setting.
The idea of using LSTM networks to decouple the actions of each time step of the prediction horizon was first explored in [11] in the context of supervised learning. In contrast, our approach employs the LSTM network as a function approximator of the decoupled Q-functions in the setting of reinforcement learning.
@@
Figure 3: A representation of the recurrent LSTM network on the left-hand side and the unrolled LSTM network on the right-hand side. Note that the LSTM is unrolled for the duration of the prediction horizon $N_{\mathrm{p}}$. Moreover, at each time step $k$, the augmented state $\chi(k)$ can go through the operation $g_{k}(\cdot)$, changing the manner in which the augmented state is presented to the LSTM network at time step $k$. This can be interpreted as a preprocessing technique to better exploit the structure of our problem.
@@
With the approximator described, the training procedure using a value-based RL method (Deep Q-Learning [34]) is discussed. Each episode of the training procedure starts with a randomly chosen initial state $\chi(k)$. Then, the RL agent repeatedly proposes a discrete action $\epsilon_{\mathrm{d}}(k)$, sends it to the environment, and collects the reward $r(k)$ and the next state $\chi(k+1)$. An episode ends when a pre-determined time limit is reached or when the RL agent suggests a discrete action that makes the optimization problem (7) infeasible. At each time step, the transitions $(\chi(k),\ \epsilon_{\mathrm{d}}(k),\ r(k),\ \chi(k+1))$ are stored in a data buffer $\mathcal{D}$, which only holds a finite number of the most recent transitions. After the end of each episode, $N_{\mathrm{batch}}$ transitions are randomly sampled from $\mathcal{D}$ and the weights of the LSTM network are updated by gradient descent on the following loss function:
$L=\frac{1}{N_{\mathrm{batch}}}\sum_{k=1}^{N_{\mathrm{batch}}}\sum_{l=0}^{N_{\mathrm{p}}-1}(y_{l}(k)-\tilde{Q}_{l}(\chi(k),\ \varepsilon_{\mathrm{d},l}(k)))^{2}$
where the temporal difference (TD) target is defined as
$y_{l}(k)=r(k)+\alpha\max_{\varepsilon}\tilde{Q}_{l}(\chi(k+1),\ \varepsilon)$
and from now on, the dependency of the decoupled Q-functions on the hidden state $h$ is omitted for ease of notation. Some alternative definitions of the TD target can be found in [36]. The training of the RL agent is deemed complete when a predetermined number of episodes or number of transitions is reached.
@@
The offline training and the online inference are described in Algorithms 1 and 2, respectively.
Algorithm 1 Offline Training
1:Initialize recurrent neural network that estimates the decoupled Q-function $\tilde{Q}(\chi,\ \epsilon;\ \theta)$ with random weights
2:Initialize databuffer $\mathcal{D}=\{\}$
3:for episode = $1,...,N_{\mathrm{episodes}}$ do
4:     Set random state $\chi(0)$
5:     for $k=1,....,N_{\mathrm{steps}}$ do
6:         for $l$ from 00 to $N_{\mathrm{p}}-1$ do
7:              $\varepsilon_{\mathrm{d},l}(k)\sim\mathrm{softmax}\ \tilde{Q}_{l}(\chi(k),\ \epsilon_{\mathrm{d},l};\ \theta)$
8:         end for
9:         Set $\epsilon_{\mathrm{d}}(k)=[\varepsilon_{\mathrm{d},0}^{\mathrm{T}}(k),\ \varepsilon_{\mathrm{d},1}^{\mathrm{T}}(k),\ \ldots,\ \varepsilon_{\mathrm{d},N_{\mathrm{p}}-1}^{\mathrm{T}}(k)]^{\mathrm{T}}$
10:         Get $\epsilon_{\mathrm{c}}(k)$ by solving the linear program (7)
11:         Set $\epsilon(k)=[\epsilon_{\mathrm{d}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{c}}^{\mathrm{T}}(k)]^{\mathrm{T}}$
12:         Get reward $r(k)=f_{\mathrm{reward}}(\chi(k),\ \epsilon(k))$
13:         Get next state $\chi(k+1)=f_{\mathrm{MDP}}(\chi(k),\ \epsilon(k))$
14:         Store $(\chi(k),\epsilon_{\mathrm{d}}(k),r(k),\chi(k+1))$ in $\mathcal{D}$
15:         Sample a minibatch of transitions $\{(\chi^{j},\epsilon_{\mathrm{d}}^{j},r^{j},\chi^{j+1})\}_{j=1}^{N_{\mathrm{batch}}}$ from $\mathcal{D}$
16:         Set $y^{j}_{l}=r^{j}$ for terminal $\chi^{j+1}$ or $y^{j}_{l}=r^{j}_{l}+\alpha\max_{\epsilon}Q_{l}(\chi^{j+1},\epsilon;\ \theta)$ for non-terminal $\chi^{j+1}$ for $l=0,...,N_{\mathrm{p}}-1$ and $j=1,...,N_{\mathrm{batch}}$
17:         Perform a gradient descent step on $\sum_{j=1}^{N_{\mathrm{batch}}}\sum_{l=0}^{N_{\mathrm{p}}-1}(y^{j}_{l}-\tilde{Q_{l}}(\chi^{j},\varepsilon_{\mathrm{d},l}^{j};\ \theta))^{2}$ to update the network weights $\theta$
18:     end for
19:end for
Algorithm 2 Online Inference
1:Measure state $\chi(k)$
2:for $l$ from 00 to $N_{\mathrm{p}}-1$ do $\triangleright$ Compute sub-actions
3:     $\varepsilon_{\mathrm{d},l}(k)=\arg\max_{\varepsilon}\ \tilde{Q}_{l}(\chi(k),\ \varepsilon)$
4:end for
5:Set = $\epsilon_{\mathrm{d}}(k)=[\varepsilon_{\mathrm{d},0}^{\mathrm{T}}(k),\ \varepsilon_{\mathrm{d},1}^{\mathrm{T}}(k),\ \ldots,\ \varepsilon_{\mathrm{d},N_{\mathrm{p}}-1}^{\mathrm{T}}(k)]^{\mathrm{T}}$
6:Get $\epsilon_{\mathrm{c}}(k)$ by solving the linear program (7)
7:return $\epsilon_{\mathrm{d}}(k)$, $\epsilon_{\mathrm{c}}(k)$